# GenAudit: Fixing Factual Errors in Language Model Outputs with Evidence

### v1

@@ Abstract
LLMs can generate factually incorrect statements even when provided access to reference documents. Such errors can be dangerous in high-stakes applications (e.g., document-grounded QA for healthcare or finance). We present GenAudit — a tool intended to assist fact-checking LLM responses for document-grounded tasks. GenAudit suggests edits to the LLM response by revising or removing claims that are not supported by the reference document, and also presents evidence from the reference for facts that do appear to have support. We train models to execute these tasks, and design an interactive interface to present suggested edits and evidence to users. Comprehensive evaluation by human raters shows that GenAudit can detect errors in 8 different LLM outputs when summarizing documents from diverse domains. To ensure that most errors are flagged by the system, we propose a method that can increase the error recall while minimizing impact on precision. We will release our tool (GenAudit) and fact-checking model for public use.
@@ 1 Introduction
LLMs can produce factually incorrect or unsubstantiated statements (Li et al., 2023a; Min et al., 2023), even when they are explicitly provided relevant context such as documents (Adams et al., 2023; Sadat et al., 2023). Incidentally, such document-grounded generation is often involved in high-stakes usage scenarios where factual correctness is (especially) paramount. For example, a doctor using an LLM to summarize a patient’s medical history (Adams et al., 2023; Kanwal and Rizzo, 2022) might make an incorrect decision if the generated summary contains errors. Manually verifying LLM outputs in such settings is therefore prudent, but also time-consuming and so undercuts the motivation for using language technologies in the first place. This motivates the need for a system that can assist users in efficiently verifying LLM output.
@@
Human evaluations also show that GenAudit can be used to verify summarization outputs in different domains, including clinical conversations, news articles and social media posts. This is despite the fact-checking model being trained only on Wikipedia data. We also evaluated its performance at factuality classification on the SummEdits benchmark (Laban et al., 2023) which consists of data from 10 different niche domains such as legal documents, scientific papers, and emails. GenAudit outperforms alternative fact-checking methods like DAE (Goyal and Durrett, 2021) and QAFactEval (Fabbri et al., 2022) and many LLMs (both open-source and proprietary), with the exceptions of GPT-4 and Gemini-pro.
GenAudit successfully identified errors in outputs from 8 different LLMs including Mistral-7B (Jiang et al., 2023), LLama2-70B (Touvron et al., 2023), Gemini-pro (Team et al., 2023) and GPT-4 (Achiam et al., 2023) in human evaluation. Observed precision ranged between $79-100\%$ while, recall ranged from $23-57\%$ for different generation models. Our human evaluation yielded a collection of $702$ summaries generated by state-of-the-art models carefully annotated with factual errors; this may be useful for future research on fact-checking LLMs.
The relative trade-off between identifying errors (recall) and making efficient use of expert time (precision) will depend on the particular use-case. We therefore introduce a decoding algorithm for fact-checking models which generate revised/fixed versions of claims, which can increase the recall of error detection with minimal cost in precision. This approach entails intervening at time-steps where the output probabilities fall below a threshold $\tau$ to select alternate decoding paths. Varying $\tau$ allows us to make more or fewer edits, effectively trading recall against precision. This approach produces a better precision-recall frontier than a baseline of randomly selecting additional words to edit to boost recall.
@@
  • We evaluate and release fine-tuned LLMs which serve as backend models for fact-checking; these perform comparably to SOTA proprietary LLMs in few-shot settings.
  • We evaluate GenAudit for fact-checking errors in summaries generated by 8 LLMs for documents from 3 domains.
  • We present and evaluate a decoding-time method that allows one to improve error detection recall (with some cost in precision).
@@ 2 Background
2.1 The USB dataset
2.2 Reducing memory requirement for training
We aim to fine-tune large models for the fact-checking task, since they are more likely to perform better and generalize to unseen domains due to internal knowledge. However, it is challenging to do so without access to large compute clusters due to memory constraints. The requirement to feed in the entire reference document to the model, which can be thousands of tokens long, further increases memory requirements.
To address the memory constraints, we use low-rank adapters with 4-bit quantization (Dettmers et al., 2023). Low rank adapters (Hu et al., 2021) reduce the memory requirement during training by reducing the number of trainable parameters, thus reducing gradient computation. To reduce the sequence length in cases where the reference document (Wikipedia article) is too long, we iteratively drop sections in it which are not relevant to any sentence in the summary (i.e. do not provide any evidence for it). We follow this process until the input sequence length is reduced to within a maximum limit, which is kept the same regardless of what model we are fine-tuning. We also use gradient accumulation and gradient checkpointing (Chen et al., 2016) to reduce memory footprint for training. For more details, please see the Appendix.
3 Experiments
Error Identification Replacements Evidence
@@ 3 Experiments
Table 3: Results from human evaluation of GenAudit predictions (using fine-tuned Flan-UL2 backend) on LLM-generated summaries of documents from different datasets.
@@ 4 Human Evaluation
Results are shown in Table 3. We use the metrics described in Section 3 for rating suggested errors and evidence generated by GenAudit. On an aggregate level—across all domains and summary generation models—GenAudit identifies erroneous words with high precision ($95.04\%$) and moderate recall ($40.37\%$). Note that achieving high recall is challenging here given the low prevalence of erroneous words ($3.97\%$). With respect to evidence extraction, we observe high precision and recall ($95.22\%$ and $90.83\%$, respectively), suggesting that most evidence sentences highlighted by the model are useful for fact-checking the given claim, and only few evidence sentences are missed (not highlighted) by the model.
@@
Using the annotations collected above, we evaluate additional models on the error detection task. Few-shot prompted GPT-4 achieves better recall than Flan-UL2, while other fine-tuned models achieve lower recall (Table 4). All models achieve lower precision than Flan-UL2. Edits suggested by GPT-4 add about $8.8\%$ more words to them on average, while other fine-tuned models add a negligible amount, reflecting the tendency of GPT-4 to make substantial changes to text.
@@
Table 4: Performance of fine-tuned models and few-shot prompted GPT-4 at identifying erroneous words in model-generated summaries, along with the percentage of summary words deleted and added by their edits.
@@ 5 Improving Recall of Error Detection
     $p_{1}p_{2}...p_{|V|}$ = ${\text{NextTokProb}}_{\mathcal{M}}(r_{1}..r_{t-1}\mid Q)$
@@
        $r^{\prime}=\arg\max_{k}(p_{k}\mid k\neq r_{t})$
@@
        compl = ${\text{Generate}}_{\mathcal{M}}(\text{prefix}\mid Q)$
        $R^{\prime}=\text{prefix}+\text{compl}$
        $N_{del},N_{add},\text{repl}$ = $\text{{DiffAtPos}}_{t}(R,R^{\prime})$
        $R=\text{prefix}+\text{repl}+r_{(t+N_{del})}...r_{|R|}$
@@
     $t=t+1$
@@
Figure 2: Variation in precision and recall of error identification by a fine-tuned Flan-UL2 model when using thresholding (Algorithm 1) versus randomly editing out additional words.
6 Binary Classification of Factuality
@@
7 Related Work
The tendency of language models to have factual errors in long-form generation was first noted in early research on machine translation (Arthur et al., 2016) and abstractive text summarization (See et al., 2017; Cao et al., 2018). Subsequent works focused on the task of classifying whether a generated summary contains any factual error or not, using trained models (Kryściński et al., 2020; Goyal and Durrett, 2021), question-answering based approaches (Fabbri et al., 2022), or prompting LLMs (Laban et al., 2023). While these works predicted factual correctness with respect to a given source document, recent works have implemented fact-checking against a large corpus (Wikipedia) by combining evidence retrieval and factuality prediction (Kamoi et al., 2023; Min et al., 2023). Our effort goes beyond binary prediction of factual correctness, by also localizing the errors in the claims and fixing them via minimal editing.
Liu et al. (2023) and Krishna et al. (2023) introduced the DeFacto and USB datasets respectively with human annotations to train and evaluate models for revising incorrect claims, and extracting evidence from a reference document. While both datasets can potentially be used to train GenAudit backend models, we used the USB dataset because of two reasons. First, USB contains comprehensive evidence labels for all facts in the claim, while DeFacto contains only evidence which justify the edits made to the claim. Second, DeFacto dataset contains mostly single-sentence summaries, while USB contains multi-sentence summaries, which are more common in practice. We extend these lines of work by contributing an interactive tool for fact-checking, and a comprehensive evaluation of the models trained on such data at fixing errors in modern LLM outputs with evidence.
Recent research has also proposed some approaches to prevent the generation of factually incorrect text from LLMs, instead of fixing the errors post-hoc. DoLa(Chuang et al., 2023) contrasts the output token logits computed at different layer outputs to promote appropriate tokens leading to more factual outputs. ITI (Li et al., 2023b) introduces biases in certain attention heads in the transformer model to promote factual outputs. While these two methods do not update model parameters,  Tian et al. (2024) improve the factuality of LLM outputs by fine-tuning them using Direct Preference Optimization (Rafailov et al., 2023).
8 Conclusion and Future Work
We introduced GenAudit, a tool to assist users in fact-checking LLM generated outputs against inputs by presenting supporting evidence and highlighting (and fixing) errors. We trained models for fact-checking tasks which rival few-shot prompting of SOTA LLMs, and designed a web-interface for users to interact with. We evaluated GenAudit for fact-checking summaries generated by 8 LLMs for documents in 3 domains. Finally, we proposed a decoding algorithm for our fact-checking model to improve the recall of error identification while minimizing the cost in precision.
An important direction for future work is to enable the use of a large corpus of reference documents instead of single reference, which could extend GenAudit to work in open-ended generation scenarios too. Although GenAudit supports real-time prediction of evidence and factuality labels as the text is edited by user, the latency is high due to the large models used. Designing smaller models for fact-checking with comparable performance but lower latency would be useful here.
Acknowledgements
Limitations
Deciding whether a span of generated text is a hallucination involves subjective judgement. While some cases clearly fall in one category (e.g. if the model invents the name of a person), others can be quite debatable. For example, while summarizing a Reddit post about a person’s bad experience, the summary mentioned them feeling left out, which was changed to disappointed by one annotator but was left as-is by the other annotator. Such decisions are motivated by the annotators’ subjective judgement of whether the inferences made in the summary are consistent with the source, leading to difference in the edits made by them. This leads to the lack of a single gold ground-truth, which makes automatic evaluation challenging.
We used the edits made to LLM-generated summaries by human annotators as gold reference while evaluating different models’ performance at error identification (Table 4). During the process, the annotators were shown suggested evidence and edits from the best-performing fine-tuned model (Flan-UL2) to evaluate its quality and to assist in the annotation process. Using the predictions from a model to assist collecting ground truth labels may lead to a benchmark that favors it and similar models. Ideally, we should conduct separate human evaluation for each model’s predictions where annotators look at the suggestions from that model before making edits. However, that would be prohibitively expensive and time-consuming, and so we conduct automatic evaluation instead by re-using the collected annotations as ground truth.
Ethical Considerations
To evaluate GenAudit, we recruited proficient proofreaders who were selected after a qualifying round, focusing on their ability to identify inaccuracies in summaries. Annotators received compensation at an average rate of \$25 USD per hour for their contributions. Annotators were provided the opportunity to discuss their concerns and questions with the authors throughout the annotation process.
The models presented in this work for fixing factual errors in LLM outputs are not perfect and some errors may not be detected by it. Hence, in critical application areas (such as clinical settings), it should be used in conjunction with a human verifier who uses GenAudit as a tool rather than as a perfect error detector. We have made this clear by transparently providing the performance of GenAudit via human evaluation in Section 4.
We verified that the different models and datasets in this work have licenses that permit research use.
Appendix A Appendix
A.1 Details on human evaluation

### v3

@@ Abstract
LLMs can generate factually incorrect statements even when provided access to reference documents. Such errors can be dangerous in high-stakes applications (e.g., document-grounded QA for healthcare or finance). We present GenAudit — a tool intended to assist fact-checking LLM responses for document-grounded tasks. GenAudit suggests edits to the LLM response by revising or removing claims that are not supported by the reference document, and also presents evidence from the reference for facts that do appear to have support. We train models to execute these tasks, and design an interactive interface to present suggested edits and evidence to users. Comprehensive evaluation by human raters shows that GenAudit can detect errors in 8 different LLM outputs when summarizing documents from diverse domains. User studies demonstrate that using GenAudit can substantially improve the performance of humans at finding errors in LLM-generated summaries. We release our tool (GenAudit) and fact-checking model for public use. [^1] https://genaudit.org
@@ 1 Introduction
LLMs can produce factually incorrect or unsubstantiated statements (Li et al., 2023; Min et al., 2023), even when they are explicitly provided relevant context such as documents (Adams et al., 2023; Sadat et al., 2023). Incidentally, such document-grounded generation is often involved in high-stakes usage scenarios where factual correctness is (especially) paramount. For example, a doctor using an LLM to summarize a patient’s medical history (Adams et al., 2023; Kanwal and Rizzo, 2022) might make an incorrect decision if the generated summary contains errors. Manually verifying LLM outputs in such settings is therefore prudent, but also time-consuming and so undercuts the motivation for using language technologies in the first place. This motivates the need for a system that can assist users in efficiently verifying LLM output.
@@
Human evaluations also showed that GenAudit can be used to verify summarization outputs in different domains, including clinical conversations, news articles and social media posts. This is despite the fact-checking model being trained only on Wikipedia data. GenAudit successfully identified errors in outputs from 8 different LLMs including Mistral-7B (Jiang et al., 2023), LLama2-70B (Touvron et al., 2023), Gemini-pro (Team et al., 2023) and GPT-4 (Achiam et al., 2023) in human evaluation. Observed precision ranged between $79-100\%$ while, recall ranged from $23-57\%$ for different generation models.
We conducted a user study to measure the impact of assistance from GenAudit on the performance of humans trying to find errors in LLM-generated summaries. Each user fact-checked a number of summaries of news articles against the corresponding references, and for half of them they were provided assistance by GenAudit. We found that the percentage of erroneous words detected by the users was $62.2\%$ when they did not use GenAudit, and it increased to $79.4\%$ when they used GenAudit. This demonstrates the ability of the tool to boost performance of humans at fact-checking.
Users fact-checking LLM outputs would not want to miss any factual errors, and may prefer that that most potential errors are highlighted for review, even if some of them are false alarms. In terms of modeling, this translates to giving a higher importance to getting better recall than precision at error-detection. We therefore introduce a decoding algorithm for fact-checking models which generate revised/fixed versions of claims, which can increase the recall of error detection while minimizing drop in precision. This approach entails intervening at time-steps where the output probabilities fall below a threshold $\tau$ to select alternate decoding paths. Varying $\tau$ allows us to suggest more or fewer edits according to user’s preference, effectively trading recall against precision. This approach produces a better precision-recall frontier compared to baselines.
@@
  • We create fine-tuned LLMs which serve as backend models for fact-checking and evaluate them comprehensively on the task of fact-checking summaries generated by 8 LLMs for documents from 3 domains.
  • We demonstrate via a user study, that using assistance from GenAudit improves performance of humans at finding errors in LLM-generated summaries.
  • We present and evaluate a custom decoding algorithm that allows one to improve error detection recall while incurring a smaller drop in precision than baselines.
@@ 2 Background
2.1 The USB dataset
2.2 Reducing memory requirement for training
We aim to fine-tune large models for the fact-checking task, since they are more likely to generalize better to unseen domains due to internal knowledge. Additionally, we need to feed in the entire reference document to the model, which can be thousands of tokens long. Both these factors increase the memory requirement for training the models, and to address this challenge, we use low-rank adapters with 4-bit quantization (Dettmers et al., 2023). Low rank adapters (Hu et al., 2021) reduce the memory requirement during training by reducing the number of trainable parameters, thus reducing gradient computation. To reduce the sequence length in cases where the reference document (Wikipedia article) is too long, we iteratively drop sections in it which are not relevant to any sentence in the summary (i.e. do not provide any evidence for it). We follow this process until the input sequence length is reduced to within a maximum limit, which is kept the same regardless of what model we are fine-tuning. We also use gradient accumulation and gradient checkpointing (Chen et al., 2016) to reduce memory footprint for training. For more details, please see the Appendix.
Figure 2: Overview of the Genaudit interface. The reference document (A) is displayed alongside the model generated output (B), allowing for easy comparison. Users can initiate the fact-checking process by clicking the fact-checking icon next to each output sentence (C), which highlights potential factual inconsistencies. Evidence markers (D) indicate supporting or contradicting evidence from the reference document. Hovering over an evidence marker highlights the relevant section in the reference (E), and clicking on it scrolls to the corresponding location in the document. Hallucinated spans in the generated output are marked in red (F), with the system providing suggestions for real-time correction
3 System Overview
The GenAudit interface presented in Figure 2, allows users to interact with both the generated output and the corresponding reference document. This interface enables users to perform the following functions aimed at enhancing the factual accuracy of generated text.
  • Review generated output: The generated text (labeled B) is displayed adjacent to the reference document (labeled A). This side-by-side layout enables an intuitive comparison between the generated text and the document it is derived from, serving as a first step in the fact-checking process.
  • Fact-Check outputs: A key feature of the interface is the ability to fact-check generated outputs with a single action. Users can fact-check any sentence in the generated text by clicking the corresponding icon next to it (labeled C), which triggers an automated verification process that cross-references information in the generated output against the reference document. Users can also click on the “Fact Check All” button to trigger fact-checking of all generated sentences.
  • Highlight Evidence: The system displays evidence markers (labeled D) next to each generated sentence. These markers link to the corresponding sentences of the reference document that provide supporting or contradicting evidence. When users hover over these evidence blobs, the relevant sentence of the reference document is highlighted (labeled E) to make the connection clear. Clicking on the evidence marker automatically scrolls the reference text to bring the supporting content into view, helping users quickly locate relevant evidence.
  • Editable Text: In cases where factual inaccuracies are detected, GenAudit offers automatic edit suggestions to correct the hallucinated spans. Users can interact with these proposed changes by reviewing the marked hallucinated spans (highlighted in red and labeled F). The system allows users to accept or reject the suggested edits, enabling real-time refinement of the generated text.
4 Experiments
@@
Error Identification Replacements Evidence Extraction
@@
Table 3: Results from human evaluation of GenAudit predictions (using fine-tuned Flan-UL2 model). Aggregate results show performance on all evaluated document-summary pairs, whereas other rows show performance (a) on summaries generated using individual LLMs, or (b) on documents from specific datasets. BaseRate represents the percentage of tokens in summaries that are hallucinations. Sufficientrepresents the percentage of summary sentences for which the evidence shown by the model was sufficient for deciding factuality.
@@ 5 Human Evaluation
Results are shown in Table 3. We use the metrics described in Section 4 for rating suggested errors and evidence generated by GenAudit. On an aggregate level—across all domains and summary generation models—GenAudit identifies erroneous words with high precision ($95.04\%$) and moderate recall ($40.37\%$). Note that achieving high recall is challenging here given the low prevalence of erroneous words ($3.97\%$). With respect to evidence extraction, we observe high precision and recall ($95.22\%$ and $90.83\%$, respectively), suggesting that most evidence sentences highlighted by the model are useful for fact-checking the given claim, and only few evidence sentences are missed (not highlighted) by the model. We have provided some samples of model-predicted errors in the Appendix, with the full collection available online [^4] https://aftersunrise.pythonanywhere.com (anonymized) .
@@
Using the annotations collected above, we evaluate additional models on the error detection task. Few-shot prompted GPT-4 achieves better recall than Flan-UL2, while other fine-tuned models achieve lower recall (Table 4). All models achieve lower precision than Flan-UL2. Edits suggested by GPT-4 add about $8.8\%$ more words to them on average, while other fine-tuned models add a negligible amount, reflecting the tendency of GPT-4 to make substantial changes to text. Finally, we also evaluate the FAVA model trained by Mishra et al. (2024) for factual error correction. We see that the model achieves the lowest recall compared to all the models evaluated, with a slightly higher precision than GPT-4.
@@
FAVA (Mishra et al., 2024) 14.18 31.34 1.80 0.43
Table 4: Performance of models fine-tuned by us on the USB dataset, few-shot prompted GPT-4, and the FAVA model (Mishra et al., 2024) at identifying erroneous words in model-generated summaries, along with the percentage of summary words deleted and added by their edits.
6 Impact on human performance at fact-checking
In the previous section, we asked humans to evaluate the correctness of the factual errors detected by GenAudit, as well as the relevance of evidence highlighted by it. While those experiments show how accurate GenAudit is, it does not reveal how helpful it would be to a human carrying out the fact-checking exercise by themselves. In this section, we measure how the fact-checking performance of a human user would change if they were given access to GenAudit.
We recruited human subjects to carry out fact-checking of summaries, and measured their performance with and without access to GenAudit predictions. In each task, users were given 2 article-summary pairs and were asked to find as many errors as they can in them. For one summary, they were shown the highlighted evidence and suggested edits from GenAudit, and for the other summary they weren’t shown the outputs of GenAudit. The rest of the interface was exactly the same. The time available to the human subject for fact-checking a summary can heavily influence their performance, and performance can perhaps increase indefinitely as more time is provided. To control for this aspect, we asked the subjects to carry out the fact-checking exercise under time constraints. We gave them 3 minutes to fact-check each summary.
The subjects were recruited using a qualification task where each candidate was given two article-summary pairs to fact-check — one with help from GenAudit and one without. The annotation exercise was attempted by $23$ candidates. We chose the top $4$ performing candidates, where their performance was measured by the F1-score of error detection.
To create the data for this experiment, we took articles from the XSUM dataset and generated summaries using GPT-3.5-turbo model. Since the rate of errors in summaries generated are quite low (ref. Table 3), most of the summaries won’t have any errors. This can be problematic, because we would like to have a relatively large number of errors to get a reliable estimate of subjects’ performance at error detection. Hence, we chose to inject errors in each summary synthetically by prompting GPT-3.5-Turbo again, following the process outlined in Laban et al. (2023) (more details are provided in the Appendix). This provided us with a larger pool of errors, as well as ground truth labels marking the erroneous spans in the summaries post-error-addition.
[…]
@@ 7 Improving Recall of Error Detection
     //  run this loop until the timestep counter reaches end of string R
     $p_{1}p_{2}...p_{|V|}$ = ${\text{NextTokProb}}_{\mathcal{M}}(r_{1}..r_{t-1}\mid Q)$run M to get probability of tokens that can be generated at timestep t}
@@
        //  if the probability of the token currently at timestep t in R is lower than tau, we replace it
        $r^{\prime}=\arg\max_{k}(p_{k}\mid k\neq r_{t})$replace it with the token with the next highest probability (call it r’)}
@@
        compl = ${\text{Generate}}_{\mathcal{M}}(\text{prefix}\mid Q)$use M to generate the full completion of the revision if we use r’ at timestep t}
        $R^{\prime}=\text{prefix}+\text{compl}$we call the new revision R’}
        $N_{del},N_{add},\text{repl}$ = $\text{{DiffAtPos}}_{t}(R,R^{\prime})$extract span-level change between R and R’ at timestep t (deletions/replacements)}
        $R=\text{prefix}+\text{repl}+r_{(t+N_{del})}...r_{|R|}$commit that change at t in R and discard any other changes observed later}
@@
     $t=t+1$advance the pointer and repeat the loop to find more spans to edit in R if they exist}
@@
We also compare using the custom decoding strategy in Algorithm 1 with simply flagging additional tokens as non-factual if their probability of generation (by the fact-checking model) falls below a variable threshold. We see that this strategy performs worse than using Algorithm 1 (Figure 3). This suggests that post-hoc usage of token probabilities from the fact-checking model does not isolate the non-factual spans as well as active intervention during the decoding process as done in Algorithm 1.
Figure 3: Variation in precision and recall of error identification by a fine-tuned Flan-UL2 model when using thresholded editing (Algorithm 1), versus editing out additional tokens either at random or by selecting the ones with low probability.
8 Related Work
With the rapid rise in capabilities of large language models (Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023), they have been widely used for assisting humans in content generation workflows such as writing summaries (Jung et al., 2023), or scientific articles (Singh et al., 2024). However, it has been known for a long time that such language models produce factually inaccurate statements, popularly known as hallucinations (Arthur et al., 2016; See et al., 2017; Cao et al., 2018). This leads to a lack of trust in the AI-generated assistance, and multiple works have noted this limitation (Qian and Wexler, 2024; Swaroop et al., 2024). To ensure factual accuracy of text, some prior works have designed intelligent user interfaces. Rony et al. (2020) introduced a tool which helps journalists fact-check claims made in debates by retrieving relevant evidence via web-search. In contrast, our work also predicts whether claims are true or false in addition to highlighting relevant evidence from reference document. Concurrent work from Fu et al. (2024) proposed a tool where claims in text are verified against structured data tables. In contrast, our system works with free-form text as reference, and does not need access to data tables whose creation can be complex and may capture only specific categories of information.
Multiple works in the natural language processing literature have introduced approaches to predict whether a generated summary contains any factual error, either using trained models (Kryściński et al., 2020; Goyal and Durrett, 2021), question-answering based approaches (Fabbri et al., 2022), or prompting LLMs (Laban et al., 2023). While these works predicted factual correctness with respect to a given source document, recent works have implemented fact-checking against a large corpus (Wikipedia) by combining evidence retrieval and factuality prediction (Kamoi et al., 2023; Min et al., 2023). Our effort goes beyond binary prediction of factual correctness, by also localizing the errors in the claims and fixing them via minimal editing. Mishra et al. (2024) also attempts to fix factual errors via editing, and we compared the performance of their released model with ours in Section 5.
Liu et al. (2023) and Krishna et al. (2023) introduced the DeFacto and USB datasets respectively with human annotations to train and evaluate models for revising incorrect claims, and extracting evidence from a reference document. While both datasets can potentially be used to train GenAudit backend models, we used the USB dataset because (1) it contains comprehensive evidence labels for all facts in the claim, and (2) it contains multi-sentence summaries, which are more common in practice. We extend these lines of work by contributing an interactive tool for fact-checking, and a comprehensive evaluation of the models trained on such data at fixing errors in modern LLM outputs with evidence.
9 Conclusion
We introduced GenAudit, a tool to assist users in fact-checking LLM generated outputs against inputs by presenting supporting evidence and highlighting (and fixing) errors. We trained models for fact-checking tasks which rival few-shot prompting of SOTA LLMs, and designed a web-interface for users to interact with. We evaluated GenAudit for fact-checking summaries generated by 8 LLMs for documents in 3 domains. We demonstrated, via a user study, that using GenAudit improves performance of humans at fact-checking summaries. Finally, we proposed a decoding algorithm for our fact-checking model to improve the recall of error identification while minimizing the cost in precision.
Acknowledgements
Appendix A Appendix
Appendix B Binary Classification of Factuality
@@
B.1 Details on human evaluation (for Section 5)